Define the following terms using equations and explain all terms used in the equation:

(a) Reynolds number (Re)

(b) Grashof’s number (Gr)

(c) Convective heat loss or gain from a body  


1
Expert's answer
2021-06-03T18:27:33-0400

a)


"Re=\frac{uL}{\nu}"


u is the flow speed

L is a characteristic linear dimension

ν is the kinematic viscosity of the fluid

b) The Grashof number is:


"Gr=\frac{g\beta (T_s-T_\infty)L^3}{\nu^2}"

where:


g is acceleration due to Earth's gravity

β is the coefficient of thermal expansion

Ts is the surface temperature

T is the bulk temperature

L is the vertical length

ν is the kinematic viscosity.

c)


"H=kA(T_s-T_a)"

where H is the loss, k is the cooling constant, A is the surface area, and Ts–Ta is the difference between skin temperature and air temperature.